OpenStudy (anonymous):

Express the radical in simplest form: radical 4/6

OpenStudy (anonymous):

\[(2\sqrt{6})/6\]

OpenStudy (anonymous):

\[\sqrt{4/6}=\sqrt{4}/\sqrt{6}=2/\sqrt{6}=2\sqrt{6}/((\sqrt{6})(\sqrt{6}))=2\sqrt{6}/6\]that should clarify any lingering steps

OpenStudy (anonymous):

You rationalize the denominator by multiplying it by \[\sqrt{6}/\sqrt{6}=1\]

OpenStudy (anonymous):

Hope that helps!